OpenStudy (deadshot):

Help? Your fishing bobber oscillates up and down from the current in the river in a harmonic motion. The bobber moves a total of 2.5 inches from its high point to its low point then returns up to the high point every 3 seconds. Write an equation modeling the motion of the bobber at its high point at time t = 0.

OpenStudy (anonymous):

the equation for a simple harmonic motion can be represented by a Sine function... that is \[f(t)=A \sin \omega t\] where A is the amplitude of oscillation omega w is the angular frequency of oscillation \[\omega=2 \pi f\] f is the frequency of oscillation

OpenStudy (anonymous):

in your problem, what do you think is A? it says, "The bobber moves a total of 2.5 inches from its high point to its low point..."

OpenStudy (deadshot):

So would A be 2.5 or 1.25?

OpenStudy (anonymous):

the amplitude is always from maximum point to minimum point...

OpenStudy (deadshot):

so then the amplitude would be 2.5

OpenStudy (anonymous):

yes...

OpenStudy (deadshot):

so that leaves us with \[f(x)= 2.5\sin wt\]

OpenStudy (anonymous):

we don't use x here, we use the time t as the abscissa...

OpenStudy (deadshot):

oh

OpenStudy (deadshot):

so then It's \[f(t)=2.5\sin wt\]

OpenStudy (anonymous):

given also the time it takes the bobber returns from low to high for time of 3 seconds, a complete period of a wave is taken from maximum to maximum (or from minimum to minimum)... so one (1) complete period of a wave is 6 seconds... knowing the period of a wave, we can determine the frequency of oscillation f... \[f=\frac{ 1 }{ period }\]

OpenStudy (anonymous):

what then would be your frequency f?

OpenStudy (deadshot):

\[f=\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

after knowing f you can now compute your angular frequency w.... then complete your equation... at any time t....

OpenStudy (anonymous):

btw according to problem, "high point at time t = 0"... meaning if t=0, f(0) = 2.5 inches... so instead of using Sine, we will use Cosine function... because the sin (0) = 0 while cos (0) =1...

OpenStudy (anonymous):

so our equation will be... \[f(t) = 2.5 \cos 2 \pi \frac{ 1 }{ 6 }t\]

OpenStudy (anonymous):

once you set t=0; f(0) = 2.5 inches.... :)

OpenStudy (deadshot):

Wait it says that the time it takes to go from high point down to low point, and then back up to high point is 3 seconds, so shouldn't the function be \[f(t)=2.5\cos2\pi \frac{ 1 }{ 3 }t\]

OpenStudy (anonymous):

oh yes i misunderstood the phrase in the sentence... it is 3 seconds not 6 seconds... sorry... :) thanks for the correction....

OpenStudy (deadshot):

No problem. Thanks for the help!